Using the accompanying Accounting Professionals data to answer the following questions. a. Find and interpret a 95​% confidence interval for the mean years of service. b. Find and interpret a 95​% confidence interval for the proportion of employees who have a graduate degree.

Employee Years of Service Graduate Degree?
1 5 N
2 10 N
3 8 Y
4 14 Y
5 8 Y
6 7 Y
7 20 N
8 6 N
9 9 Y
10 19 Y
11 10 N
12 28 N
13 9 N
14 8 N
15 8 Y
16 5 N
17 31 N
18 9 N
19 8 N
20 26 Y
Answer #1

Using Excel

a)

mean 12.4000
sd 7.9498
n 20

95% confidence interval

(Xbar +- t * s/sqrt(n))

df =n-1 = 19

alpha = 0.05

t = t.inv.2t(alpha,df)

t 2.093
t * sd/sqrt(n) 3.721
lower 8.6794
upper 16.1206

b)

Employee Years of Service Graduate Degree?
3 8 Y
4 14 Y
5 8 Y
6 7 Y
9 9 Y
10 19 Y
15 8 Y
20 26 Y

X = 8

n = 20

(p^ +- z *sqrt(p^(1-p^)/n)

p^ = X/n

z = 1.96 for 95%

p^ 0.400

E = z *sqrt(p^(1-p^)/n)

2-tailed 1.9600
E 0.2147
lower 0.1853
upper 0.6147